### v3

# Dynamic Multi-Hop HARQ (DMH-HARQ): Reliable and Open Latency-Constrained Wireless Transport Network

Abstract
The extreme requirements for high reliability and low latency in the upcoming Sixth Generation (6G) wireless networks are challenging the design of multi-hop wireless transport networks. Inspired by the advent of the virtualization concept in the wireless networks design and openness paradigm as fostered by the O-RAN Alliance, we target a revolutionary resource allocation scheme to improve the overall transmission efficiency.
In this paper, we investigate the problem of multi-hop decode-and-forward (DF) relaying in the finite blocklength (FBL) regime, and propose a DMH-HARQ scheme, which maximizes the end-to-end (E2E) communication reliability in the wireless transport network. We also propose an integer dynamic programming (DP) algorithm to efficiently solve the optimal DMH-HARQ strategy. Constrained within a certain time frame to accomplish E2E transmission, our proposed approach is proven to outperform the conventional listening-based cooperative ARQ, as well as any static HARQ strategy, regarding the E2E reliability. It is applicable without dependence on special delay constraint, and is particularly competitive for long-distance transport network with many hops.
Index Terms:
6G, URLLC, HARQ, openness, relay, FBL, dynamic programming
I Introduction
Wireless systems are conceived as a reliable and efficient means to communicate. In particular, with the last generation of wireless networks (5G) unprecedented performance have been achieved while preserving privacy and reliability. In this context, the utilization of relay nodes in wireless communications has proven to be an effective method for combating unexpected network dynamics, such as channel fading, and, in turn, extending coverage [1, 2]. However, with the emergence of novel mobile applications and their increasing demand for exceptional reliability and minimal latency, namely ultra-reliable low-latency communications (URLLC), the feasibility of employing relaying networks in such scenarios has become a challenge [3]. The usage of short codes to meet stringent latency requirements not only significantly reduces the benefits of relaying due to higher error rates [4], but also conventional automatic repeat request (ARQ) solutions struggle to effectively address the diminished link-level reliability as they typically do with longer codes [5].
In parallel, the next generation of wireless networks will power a revolution: network virtualization with a specific focus on radio access networks, a.k.a. virtualized radio access networks (vRAN) [6]. It will hold the promise of substantial cost savings in both capital and operating expenses, while also offering high flexibility and openness to foster innovation and competition. This transformative technology serves as the ultimate milestone as we progress towards the realm of 6G. Nevertheless, extending virtualization to the radio access domain presents a range of challenges that have garnered attention from various initiatives and organizations. Among those, the Open-Radio Access Network (O-RAN) alliance has emerged as a particularly influential force, gaining significant traction and support [7]. O-RAN, a carrier-led initiative, aims to define the next generation of (v)RANs for multi-vendor deployments with a primary objective to disrupt the vRAN ecosystem by dismantling vendor lock-ins and creating an open market that has traditionally been dominated by a small group of players: If O-RAN achieves its goals, it has the potential to unleash an unparalleled level of innovation by reducing the barriers to entry for new competitors [8].
However, the practical deployment of O-RAN in future Sixth Generation (6G) system is facing various emerging challenges, among which there is the essential introduction of wireless transport network to replace the fiber solution. The demand for wireless transport network in 6G is twofold. On the one hand, to support higher throughput and enable accurate positioning, new spectrum at higher frequencies ranging from millimeter-wave (mmWave) to terahertz (THz) will certainly be exploited. The dependence on line-of-sight (LoS) link of these frequencies, however, is implying limited coverage of every individual base station (BS), and inevitably leading to an ultra-dense deployment of small cells [9]. The conventional fiber technology, unfortunately, fails to serve as a reasonably cost-efficient solution for transport networks in this context [9, 10]. On the other hand, aiming at a ubiquitous three-dimensional (3D) coverage, 6G is supposed to deploy non-terrestrial infrastructures such as unmanned aerial vehicles (UAVs), high altitude platforms (HAPs), and satellites, as its serving nodes [11]. Wireless transport network plays here the role of a fundamental enabling technology for the non-terrestrial network (NTN).
In the 6G O-RAN transport network, multi-hop relay mechanisms are integral to the system architecture. User data and control signaling, irrespective of their origin – whether from user equipment (UE) in the uplink (UL) or destined for UE in the downlink (DL)– undergo a sequential relay process through the network’s architecture. The relaying can be executed by a series of network components, including the radio hardware unit (RU), the distributed unit (DU), the central unit (CU), and eventually multiple relay nodes in the case of a long-distance backhaul. Depending on the network functions involved in the communication session and their placement in the network, either part or all of these components can be involved in the relay chain.
Implementing this multi-hop transport network over wireless channels, especially in the context of 6G O-RAN, is challenging conventional radio interface protocols in two aspects. First, compared to classical fiber solutions, wireless technologies are not only suffering from higher propagation loss, exposed to stronger noise and interference, but also more strictly regulated regarding the transmission power. These facts jointly lead to a low signal-to-interference-and-noise ratio (SINR), which challenges the reliability of the radio link, especially when the latency constraint is stringent. Concerning the ambition of 6G for ubiquitous coverage, it will not be a rare case to have long relay chains with many hops in the transport network, which further escalates the conflict between the 6G demands for reliability and latency. To fulfill these requirements, novel methods will be essential to push the transmission performance beyond the state of the art. Second, supporting flexible functional splitting, the O-RAN architecture is also implying a dynamic setup of the relay chain upon the specific choice among different splitting options. Meanwhile, the hybrid 6G architecture consisting of both terrestrial and non-terrestrial infrastructure increases the diversity of single-hop channel features in different deployment scenarios. The high dynamics and diversity of the multi-hop transport network are challenging conventional radio protocols based on fixed radio frame design and static resource scheduling.
Motivated by the need to enhance openness and reliability to end-to-end (E2E) links while fulfilling strict time constraints imposed by URLLC applications in a multi-hop transport netork, we draw inspiration from the concept of dynamically reallocating the remaining radio resources between repeats and successor hops based on instantaneous transmission results. Therefore, in this paper we pioneer the design of a novel dynamic hybrid ARQ (HARQ) scheme for multi-hop relaying in the finite blocklength (FBL) regime. We conduct a comprehensive analysis of its performance and propose a dynamic programming (DP) algorithm for optimizing its schedule. By exploring this approach, we aim to address the challenges faced by traditional relaying networks and improve the overall reliability of the O-RAN-inspired wireless transport network. Our research focuses on achieving enhanced E2E reliability through dynamic resource allocation, taking into account the specific requirements and constraints of modern mobile systems operating in high-demand scenarios.
The remainder of the paper is organized as follows: We begin with Section II to provide a brief review to selected literature in related field. Section III follows to describe and formulate the investigated problem, which is then analyzed in Section IV and solved in Section V. The proposed methods are numerically evaluated in Section VI, followed by some further discussions in Section VII. To the end, we close the paper with our conclusion and outlooks in Section VIII.
II Related Work
The problem of multi-hop relay has been attracting great interest in the wireless field since long, for it has a rich potential to offer extended cell coverage and improved channel capacity [1], particularly in the O-RAN architecture [12, 13]. Generally, there are two families of relay technologies, namely decode-and-forward (DF) and amplify-and-forward (AF). Albeit the disadvantage of higher implementation complexity, DF comprehensively outperforms AF in latency-reliability performance [14], and is therefore preferred by many researchers. Literature has investigated multi-hop DF from various perspectives, including modulation scheme [15], selective relaying route [16], and relay station placement [17], and ARQ strategy [18].
Over the past few years, the emerging wireless use scenario of URLLCin O-RAN is challenging off-the-shelf multi-hop relay. The extreme requirement of low E2E latency in O-RAN URLLC is leading to even more stringent time constraint over each individual hop [19, 20], which is barely achievable with conventional technologies. To meet such a stringent latency requirement, URLLC transmissions are likely carried out with short packet, i.e., the codewords cannot considered as infinite long. With such so-called FBL codes, many theories and propositions in classical information theory fail to apply since the Shannon limit is no more asymptotically achievable. To accurately characterize the relationship between achievable data rate and reliability in the FBL regime, Polyanskiy et al. derived in [21] a closed-form expression for the single-hop transmission in additive white Gaussian noise (AWGN) channels. Later on, this FBL characterization was generalized, extended into Gilbert-Elliot channels [22] and quasi-static flat-fading channels [23]. Recently, the FBL model in random access channel has also been analyzed [24]. On the basis of those results, the performance FBL communication systems has been investigated in context of various wireless networking technologies, such as multi-input multi-output (MIMO) [25], orthogonal frequency-division multiple access (OFDMA) [26], non-orthogonal multiple access (NOMA) [27] and multi-access edge computing (MEC) [28].
Specifically for FBL relay networks, the system performance are generally investigated in a deterministic manner in the most existing works [4, 29, 30, 31, 32]. In particular, the authors of [4] studied the performance bound of two-hop FBL relay system with perfect channel state information (CSI), which was later extended to the scenarios with only partial CSI available [29]. Moreover, the authors of [30] investigated a potential use case where UAV behaves as the relay between source and destination to improve the system performance. Energy harvesting was introduced in [31] where the relay solely relies on the harvested energy to forward the information. The authors of [32] studied the impact of imperfect successive interference cancellation (SIC) in NOMA schemes on relay system that causes by the FBL transmission error.
Another interesting FBL use case is the closed-loop communication, which is not only common in MECscenarios but also strongly related to FBL relay networks, because it behaves in a similar way as a two-hop relay chain does. In a previous work of ours we proposed for this use case the so-called Closed-Loop ARQ (CLARQ) algorithm [33], which dynamically re-allocates blocklengths between uplink retransmission and downlink reception, so as to minimize the closed-loop error rate. Moreover, in [34] we analyzed the optimal one-shot transmission scheme, which is a special case of CLARQ, under both constraints of latency and energy consumption.
III Problem Setup
[…]
III-A Multi-Hop DF and DMH-HARQ
Now consider a multi-hop relaying chain with $I+1$ nodes over $I$ hops, each applying the DF strategy based on simple HARQ without combining. We assume that all channels of the $I$ hops are i.i.d., undergoing Gaussian noise and Rician fading. We assume that every decoding process, together the associated feedback of acknowledgement (ACK) or negative acknowledgement (NACK) to the predecessor node for the HARQ procedure, makes the same delay of $T_{\mathrm{l}}$. The overall transmission to a maximal E2E delay is limited to $T_{\mathrm{max}}$. For simplification of the analysis, we consider for every node the same symbol rate $f_{\mathrm{s}}$, modulation order $M$ (and therefore the bit-time length $T_{\mathrm{b}}$=$\frac{1}{Mf_{\mathrm{s}}}$), transmission power $P$, and noise power $N$. The transmitting node of each hop is capable of flexibly adjust its channel coding rate in every (re-)transmission attempt independently from others. We consider in this study a block fading scenario, where the channel $h_{i}$ of each hop $i\in\mathcal{I}=\{1,2,\dots I\}$ remains consistent over the entire time frame of $T_{\mathrm{max}}$. Each transmitting node of hop $i$ possesses the perfect CSI of $h_{i}$, while knowing only its position in the relaying chain and the statistical CSI of all other hops.
Here we consider a problem of dynamic radio resource allocation among different HARQ slots over the $I$-hop DF chain. On every hop, the transmitting node can individually adjust its coding rate for the next (re)transmission. Since the modulation order and symbol rate are fixed, this will determine the time span of the transmission slot, which cannot exceed the remaining time of the entire transmission frame. If the message is successfully decoded, the successor node will take the turn and further forward the message, unless it is the sink node, i.e. the $(I+1)^{\mathrm{th}}$. If the decoding fails, the receiving node will trigger the HARQ procedure by sending back a NACK to the transmitting node, which then keeps re-transmitting the message until either a successful reception or a timeout, i.e. when the remaining time falls insufficient to support any further effective transmission. In every individual repeat, the node is able to readjust its coding rate. We consider a timer embedded in the header of every message, so that every node is aware of the time remaining from the $T_{\mathrm{max}}$ frame. We name this proposed approach as Dynamic Multi-Hop HARQ (DMH-HARQ), which is briefly illustrated in Fig. 1.
Figure 1: Illustration of a DMH-HARQ schedule for 4-hop DF chain, where the remaining (reserved) blocklength of a frame is dynamically assigned regarding the decoding result: either for next hops upon success, or for retransmission on the same hop upon failure.
III-B Binary Schedule Tree
From Fig. 1 we see that the complete schedule of DMH-HARQ has a structure of complete binary tree $\mathfrak{T}$, which can be recursively defined as a 3-tuple $\mathfrak{T}=(L(\mathfrak{T}),S(\mathfrak{T}),R(\mathfrak{T}))$, where $S(\mathfrak{T})$ is a singleton set containing the root node of $\mathfrak{T}$, $L(\mathfrak{T})$ and $R(\mathfrak{T})$ are the binary trees rooted on the left and right children of the root node, respectively:
$L(\mathfrak{T})=\begin{cases}\emptyset&S(\mathfrak{T})=\emptyset,\\ \left(L(L(\mathfrak{T})),S(L(\mathfrak{T})),R(L(\mathfrak{T}))\right)&\text{o.w.},\end{cases}$ (1)
$R(\mathfrak{T})=\begin{cases}\emptyset&S(\mathfrak{T})=\emptyset,\\ \left(L(R(\mathfrak{T})),S(R(\mathfrak{T})),R(R(\mathfrak{T}))\right)&\text{o.w.}\end{cases}$ (2)
Each node $s$ in the tree representing a certain system state, which can be defined as a three tuple $s=(t_{s},i_{s},k_{s},\tau_{s})$ where $t_{s}\in\mathbb{R}^{+}$ is the remaining time, $i_{s}\in\mathbb{N}$ is the remaining hops to transmit the message, $k_{s}\in\mathbb{N}$ is the number of failed HARQ attempts on the current hop, and $\tau_{s}$ the accumulated time spent on these $k_{s}$ unsuccessful transmissions (excluding the feedback losses). Now let $n_{s}\triangleq n(t_{s},i_{s},k_{s},\tau_{s})\in\mathbb{N}$ denote the the blocklength to be allocated, according to the policy $n(\cdot)$, to the current HARQ attempt on the current hop, i.e. the $(k_{s}+1)$th HARQ attempt on the $i_{s}$th last hop, when there is $t_{s}$ time remaining. Thus, let $\mathfrak{T}_{s}$ denote a schedule tree with $S(\mathfrak{T}_{s})=\{s\}$ and $\mathcal{S}$ the space of all feasible states, for all $s\in\mathcal{S}$, on the one hand we have for the left branch
$\begin{split}&S(L(\mathfrak{T}_{s}))\\ =&\begin{cases}\emptyset&t_{s}<T_{\mathrm{min}}^{(i_{s})},\\ \emptyset&i_{s}=0,\\ \left\{(t_{s}-n_{s}T_{\mathrm{b}}-T_{\mathrm{l}},i_{s},k_{s}+1,\tau_{s}+n_{s}T_{\mathrm{b}})\right\}&\text{o.w.}\end{cases}\end{split}$ (3)
Here, $T_{\mathrm{min}}^{(i)}=\sum\limits_{j=1}^{i}\left(n_{\mathrm{min},j}T_{\mathrm{b}}+T_{\mathrm{l}}\right)$ where $n_{\mathrm{min},j}$ is the minimal blocklength of codeword on the $j$th last hop (which is a function of the signal-to-noise ratio (SNR) $\gamma_{i}$ over the $j$th last hop). Therefore, in the first case of Eq. (3), $t_{s}<T_{\mathrm{min}}^{(i)}$ denotes the condition that a remaining time $t_{s}$ is insufficient to forward the message over the last $i$ hops even when no error occurs (i.e. with only one transmission attempt on each hop), under which the DMH-HARQ procedure is terminated without success. The second case a successful completion of the DMH-HARQ over the entire chain (as there is $i=0$ hop remaining); the third case stands for the consequence when the current HARQ attempt fails with sufficient remaining time, i.e. the message is re-transmitted over the current hop for the $(k_{s}+1)$th time.
On the other hand, for the right branch we have
$S(R(\mathfrak{T}_{s}))=\begin{cases}\emptyset&t_{s}<T_{\mathrm{min}}^{(i_{s})},\\ \emptyset&i_{s}=0,\\ \left\{(t_{s}-n_{s}T_{\mathrm{b}}-T_{\mathrm{l}},i_{s}-1,0,0)\right\}&\text{o.w.}\end{cases}$ (4)
Again, the first case implies to insufficient remaining time and the second case describes a successful completion of DMH-HARQ. The third case stands for a successful DF over an intermediate hop, when the DMH-HARQ progresses to the next hop, i.e. the $(i_{s}-1)$th last hop.
III-C Optimization Problem
To measure the effectiveness of a DMH-HARQ schedule, we define a utility $\xi(\mathfrak{T}_{s})$ as the chance of successfully forwarding the message over the next $i_{s}$ hops within a time limit of $t_{s}$, then from the recursive structure (12) of $\mathfrak{T}_{s}$, we have
$\xi(\mathfrak{T}_{s})=\begin{cases}0&\mathfrak{T}_{s}=\emptyset,\\ 1&i_{s}=0,\\ \varepsilon_{s}\xi(L(\mathfrak{T}_{s}))+(1-\varepsilon_{s})\xi(R(\mathfrak{T}_{s}))&\text{o.w.},\end{cases}$ (5)
where $\varepsilon_{s}$ is the packet error rate (PER) of the $(k_{s}+1)$th HARQ attempt on the $i_{s}$th last hop with codeword blocklength $n_{s}$. Since $\varepsilon_{s}\in[0,1]$, we know that $\xi(\mathfrak{T}_{s})$ is bounded between $[0,1]$ for any $\mathfrak{T}_{s}$.
[…]
$n_{\mathrm{min},i_{s}}=\min\{n\in\mathbb{N}^{+}~{}|~{}\varepsilon_{s}\leqslant\varepsilon_{\mathrm{max}}\}$ (6)
IV Analyses: Type I HARQ without Combining
Before diving deep into the optimal HARQ scheduling problem, we shall study the error model of each HARQ attempt. For a delay-critical scenario with stringent E2E latency constraint, such as URLLC, we consider the use of short code, where the blocklength of every codeword is limited. Therefore, the PER of every HARQ attempt shall be analyzed w.r.t. the finite blocklength information theory, and the performance highly depends on the selected coding and HARQ schemes.
The choice of combining technique can have crucial impact in the performance of HARQ protocols. It is well known that among the different types of HARQ, Type I HARQ without combining provides the worst error performance with the simplest implementation, while the Type II HARQ-incremental redundancy (IR) offers the best error performance with the highest implementation complexity. On the one hand, applying Type II HARQ-IR to our DMH-HARQ protocol will certainly grant it an optimal performance. Unfortunately, however, the generic error model will therewith become analytically intractable, due to the dual inconsistencies – regarding both the redundancy rate and the frame length – over individual transmission attempts. While numerical simulations can be conducted to obtain the empirical model dedicated to any certain channel coding scheme, such models lack generality and therefore have limited theoretic contribution. On the other hand, the Type I HARQ without combining, known as the worst scheme regarding error performance, has a rich value in analysis since it outlines a tight upper-bound for PER of all HARQ solutions. Therefore, in this study we focus on the case of Type I HARQ without combining, where the receiver does not exploit any information of the unsuccessfully decoded packets.
IV-A Error Model, Constraints, and Approximations
Regarding Type I HARQ without combining, according to [21], the PER of an arbitrary HARQ attempt with blocklength $m$ on the $i^{\mathrm{th}}$ hop is
$\epsilon_{i}(m)=Q\left(\sqrt{\frac{m}{V_{i}}}\left(\mathcal{C}_{i}-\frac{d}{m}\right)\ln 2\right),$ (7)
where for AWGN channels $V_{i}=1-\frac{1}{(1+\gamma_{i})^{2}}$ with $\gamma_{i}=\frac{Pg_{i}}{N}=\frac{P\overline{g}}{NL_{\text{f},i}}.$ denoting the SNR on the $i^{\mathrm{th}}$ hop. Here, $g_{i}$ is the channel gain of the $i^{\mathrm{th}}$ hop, which is determined by the mean channel gain $\overline{g}$ and the random fading loss $L_{\text{f},i}$. Thus, for all $s=(t_{s},i_{s},k_{s},\tau_{s})\in\mathcal{S}$, with the blocklength allocation policy $n_{s}$, we have
$\varepsilon_{s}=\epsilon_{I+1-i_{s}}(n_{s})=Q\left(\sqrt{\frac{n_{s}}{V_{I+1-i_{s}}}}\left(\mathcal{C}_{I+1-i_{s}}-\frac{d}{n_{s}}\right)\ln 2\right),$ (8)
which has no dependency on $t_{s}$, $k_{s}$ or $\tau_{s}$ except for the potential dependency of $n_{s}$ on them.
Following the common routine in FBL study, we can extend the space of blocklength from integer to real numbers, so as to allow non-integer values of $n_{s}$ for the convenience of analysis (approximate integer solutions can be then obtained by rounding the real-valued optima). Thus, the original problem (III-C) can be replaced by {maxi} n:SR^+ξ(T_s_0)
Recalling (6), the lower-bound $n_{\mathrm{min},i_{s}}$ in constraint (III-C) can be obtained from:
$Q\left(\sqrt{\frac{n_{\mathrm{min},i_{s}}}{V_{I+1-i_{s}}}}\left(\mathcal{C}_{I+1-i_{s}}-\frac{d}{n_{\mathrm{min},i_{s}}}\right)\ln 2\right)=\varepsilon_{\mathrm{max}},$ (9)
which depends on $s$ only regarding $i_{s}$. Meanwhile, the upper-bound $n_{\mathrm{max},s}\triangleq\frac{t_{s}-T_{\mathrm{min}}^{(i_{s}-1)}}{Mf_{\mathrm{s}}}$ is determined by both $t_{s}$ and all $n_{\mathrm{min},j}$ that $11\leqslant j\leqslant i_{s}-$ 1, i.e., by $t_{s}$ and $i_{s}$.
IV-B Multi-Hop Link Reliability with Perfect Global CSI
Based on the error model above, we start our analysis to the optimization problem (III-C) with a simplified case where perfect CSI of all hops $i\in\mathcal{I}$ is available at every node. For the convenience of notation, in the analyses below we let
$\xi_{\mathrm{max}}(s)\triangleq\max\limits_{n:\mathcal{S}\to\mathbb{R}^{+}}\xi(\mathfrak{T}_{s}),\qquad\forall s\in\mathcal{S},$ (10)
and the corresponding optimal blocklength allocation policy
$n_{\mathrm{opt}}(s)\triangleq\arg\max\limits_{n:\mathcal{S}\to\mathbb{R}^{+}}\xi(\mathfrak{T}_{s}),\qquad\forall s\in\mathcal{S}.$ (11)
First, recall the remarks we made earlier in Section IV-A, that given a certain $s$, both the PER $\varepsilon_{s}$ and the constraints to $n_{s}$ are independent from $k_{s}$ or $\tau_{s}$, but determined by $t_{s}$ and $i_{s}$. Thus, we can assert that
Lemma 1.
With Type I HARQ without combining, $n_{\mathrm{opt}}(s)$ is independent from $k_{s}$ or $\tau_{s}$, but determined by $t_{s}$ and $i_{s}$, i.e. $n_{\mathrm{opt}}(s)=n^{\mathrm{opt}}_{i_{s}}(t_{s})$.
The proof of Lemma 1 is trivial and therefore omitted.
For convenience we let $s_{\mathrm{l}}(\mathfrak{T}_{s})$ denote the root state of $L(\mathfrak{T}_{s})$, and similarly $s_{\mathrm{r}}(\mathfrak{T}_{s})$ the root state of $R(\mathfrak{T}_{s})$. Thus, from the recursive structure Eq. (5) of the utility $\xi(\mathfrak{T}_{s})$, for all $s$ that $\mathfrak{T}_{s}\neq\emptyset$ and $i_{s}\neq 0$, we can derive
$\begin{split}&\xi_{\mathrm{max}}(s)\\ =&\max\limits_{n_{s}}\left\{\varepsilon_{s}\xi_{\mathrm{max}}(s_{\mathrm{l}}(\mathfrak{T}_{s}))+(1-\varepsilon_{s})\xi_{\mathrm{max}}(s_{\mathrm{r}}(\mathfrak{T}_{s}))\right\}\\ =&\max\limits_{n_{s}}\left\{\varepsilon_{s}\max\limits_{n_{S(L(\mathfrak{T}_{s}))}}\xi(L(\mathfrak{T}_{s}))+(1-\varepsilon_{s})\max\limits_{n_{S(R(\mathfrak{T}_{s}))}}\xi(R(\mathfrak{T}_{s}))\right\}\\ =&\dots\end{split}$ (12)
which forms a Bellman equation. Under the error model (8), we have the following conclusions:
Lemma 2.
With Type I HARQ without combining, for all $s\in\mathcal{S}$ that $i_{s}=1$ (i.e. on the last hop), the optimal blocklength allocation is always a one-shot transmission $n_{\mathrm{opt}}(s)=\frac{t_{s}-T_{\mathrm{l}}}{T_{\mathrm{b}}}$ that maximizes the utility to $\xi_{\mathrm{max}}(s)=1-\epsilon_{I}\left(\frac{t_{s}-T_{\mathrm{l}}}{T_{\mathrm{b}}}\right)$.
Theorem 1.
As long as $t_{s}\geqslant T_{\mathrm{min}}^{(i_{s}-1)}$ and $i_{s}\geqslant 1$, $\xi_{\mathrm{max}}(s)$ is a continuous and strictly monotone increasing function of $t_{s}$.
The proofs are provided in Appendices A and B, respectively.
IV-C Multi-Hop Link Reliability with Imperfect CSI
As we have discussed in Section III, in the multi-hop relaying scenario, it shall be considered that each node $i$ has only a perfect CSI of the $i^{\mathrm{th}}$ hop. For all the successor hops $\{i+1,i+2,\dots I\}$, it possesses only the statistical CSI, i.e. the probability density function (PDF) $f_{G_{i}}$ of their channel gains $g_{i}$, which can be reasonably assumed consistent and identical for every hop $i\in\mathcal{I}$, i.e. $f_{G_{i}}\equiv f_{G}$ for all $i\in\mathcal{I}$. Thus, in an arbitrary state $s\in\mathcal{S}$ where the node $I+1-i_{s}$ is forwarding the message, the node cannot accurately estimate $\varepsilon_{s^{\prime}}$ for any state $s^{\prime}$ that $i_{s^{\prime}}<i_{s}$ upon the blocklength allocation $n_{s^{\prime}}$ anymore, but only the expectation upon the random channel gain:
$\overline{\varepsilon}_{s^{\prime}}=\int\limits_{0}^{+\infty}\varepsilon_{s^{\prime}}f_{G}(g_{i_{s}^{\prime}})\text{d}g_{i_{s^{\prime}}}.$ (13)
Note that for any specific $\gamma_{i_{s^{\prime}}}\geqslant 0$, $\varepsilon_{s^{\prime}}$ is convex and monotone decreasing w.r.t. $n_{s^{\prime}}$. Thus, as a linear combination of various samples of $\varepsilon_{s^{\prime}}$ over $\gamma_{i_{s^{\prime}}}\in[0,+\infty)$, $\overline{\varepsilon}_{s^{\prime}}$ is also convex and monotone decreasing w.r.t. $n_{s^{\prime}}$, so that Theorem 1 remains applicable even when only statistical CSI of for all successor hops is available at each node.
V Algorithm Design
V-A Optimal DMH-HARQ: Integer Dynamic Programming
From Lemma 1 we know that when applying Type I HARQ without combining, the DMH-HARQ scheduling is a typical Markov decision process (MDP), where the expected utility $\xi_{s}$ from any certain status $s$ is independent from its historical states or actions, but only relying on the remaining time $t_{s}$ and the blocklength allocation policy $n_{s}$. Such kind of problems can be generally solved by recursive approaches that gradually evaluate and improve the decision policy by traversing the entire feasible region of states, typical examples are policy iteration and value iteration. However, such approaches are challenged by two issues when dealing with Eq. (12). First, (8) provides no closed-form analytical solution $n_{s}$ for an arbitrary $\varepsilon_{s}$. Second, with $n_{s}\in\mathbb{R}^{+}$, both the feasible region $\mathcal{S}$ and the set of possible actions becomes infinite, making it impossible to traverse the solution space within limited time.
{strip}
1
2function OptSchedule($s,g_{i_{s}},f_{G},\varepsilon_{\mathrm{max}},T_{\mathrm{b}},T_{\mathrm{l}}$):
       $\mathcal{D}_{\mathrm{local}}\leftarrow\mathrm{Dict}([~{}]),\mathcal{D}_{\mathrm{succ}}\leftarrow\mathrm{Dict}([~{}])$
        // Empty dictionaries.
       Best!($t_{s},i_{s},g_{i_{s}},f_{G},\varepsilon_{\mathrm{max}},\mathcal{D}_{\mathrm{local}},\mathcal{D}_{\mathrm{succ}}$)
        // Solve the optimal blocklength allocation policy.
       $\mathfrak{T}_{s}\leftarrow$ScheduleTreeGen($t_{s},i_{s},0,0,\mathcal{D}_{\mathrm{local}},T_{\mathrm{b}},T_{\mathrm{l}}$)
        // Generate the binary schedule tree.
3
4
5return $\mathfrak{T}_{s}$
6
7
8function Best!($t_{s},i_{s},g_{i_{s}},f_{G},\varepsilon_{\mathrm{max}},\mathcal{D}_{\mathrm{local}},\mathcal{D}_{\mathrm{succ}}$):
       $n_{\mathrm{min},i_{s}}\leftarrow\min\{n:\varepsilon_{s}(n)\leqslant\varepsilon_{\mathrm{max}}\}$
        // W.r.t. Eq. (8).
       $\overline{n}_{\mathrm{min}}\leftarrow\min\{n:\overline{\varepsilon}(n)\leqslant\varepsilon_{\mathrm{max}}\}$
        // W.r.t. Eq. (13).
9       if $\mathcal{D}_{\mathrm{local}}.\mathrm{find}\left([t_{s},i_{s}]\right)\neq\emptyset$ then
             $[\xi_{\mathrm{opt}},n_{\mathrm{opt}}]\leftarrow\mathcal{D}_{\mathrm{local}}.\mathrm{find}\left([t_{s},i_{s}]\right)$
              // Already solved, directly read from the look-up table (LUT).
10
11       else if $t_{s}<\left[(i_{s}-1)\overline{n}_{\mathrm{min}}+n_{\mathrm{min},i_{s}}\right]T_{\mathrm{b}}+i_{s}T_{\mathrm{l}}$ then
             $[\xi_{\mathrm{opt}},n_{\mathrm{opt}}]\leftarrow[0,0]$
              // Insufficient radio resource left, DMH-HARQ terminated.
             $\mathcal{D}_{\mathrm{local}}.\mathrm{append}\left([t_{s},i_{s}]\rightarrow[\xi_{\mathrm{opt}},n_{\mathrm{opt}}]\right)$
              // Append the new result to LUT.
12
13       else if $i_{s}==1$ then
             $\left[\xi_{\mathrm{opt}},n_{\mathrm{opt}}\right]\leftarrow\left[\varepsilon_{s}(n_{\mathrm{opt}}),\left\lfloor\frac{t_{s}-T_{\mathrm{l}}}{T_{\mathrm{b}}}\right\rfloor\right]$
              // Last hop, one-shot as optimum.
14             $\mathcal{D}_{\mathrm{local}}.\mathrm{append}\left([t_{s},i_{s}]\rightarrow[\xi_{\mathrm{opt}},n_{\mathrm{opt}}]\right)$
15       else
16             $n_{\mathrm{opt}}\leftarrow\arg\max\limits_{n}\{\varepsilon_{s}(n)\textnormal{{Best!}}(t_{s}-nT_{\mathrm{b}}-T_{\mathrm{l}},i_{s},g_{i_{s}},f_{G},\varepsilon_{\mathrm{max}},\mathcal{D}_{\mathrm{local}},\mathcal{D}_{\mathrm{succ}})$
17             ${}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+[1-\varepsilon_{s}(n)]\textnormal{{BestSucc!}}(t_{s}-nT_{\mathrm{b}}-T_{\mathrm{l}},i_{s}-1,f_{G},\varepsilon_{\mathrm{max}},\mathcal{D}_{\mathrm{succ}})\}$
18             $\xi_{\mathrm{opt}}\leftarrow\varepsilon_{s}(n_{\mathrm{opt}})\textnormal{{Best!}}(t_{s}-n_{\mathrm{opt}}T_{\mathrm{b}}-T_{\mathrm{l}},i_{s},g_{i_{s}},f_{G},\varepsilon_{\mathrm{max}},\mathcal{D}_{\mathrm{local}},\mathcal{D}_{\mathrm{succ}})+$
19             ${}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}[1-\varepsilon_{s}(n_{\mathrm{opt}})]\textnormal{{BestSucc!}}(t_{s}-n_{\mathrm{opt}}T_{\mathrm{b}}-T_{\mathrm{l}},i_{s}-1,f_{G},\varepsilon_{\mathrm{max}},\mathcal{D}_{\mathrm{succ}})$
20             $\mathcal{D}_{\mathrm{local}}.\mathrm{append}\left([t_{s},i_{s}]\rightarrow[\xi_{\mathrm{opt}},n_{\mathrm{opt}}]\right)$
21       end if
22
23
24return $\xi_{\mathrm{opt}}$
25
26
27function BestSucc!($t_{s},i_{s},f_{G},\varepsilon_{\mathrm{max}},\mathcal{D}$):
28       $\overline{n}_{\mathrm{min}}\leftarrow\min\{n:\overline{\varepsilon}(n)\leqslant\varepsilon_{\mathrm{max}}\}$
29       if $\mathcal{D}.\mathrm{find}\left([t_{s},i_{s}]\right)\neq\emptyset$ then
30             $[\xi_{\mathrm{opt}},n_{\mathrm{opt}}]\leftarrow\mathcal{D}.\mathrm{find}\left([t_{s},i_{s}]\right)$
31       else if $t<i(\overline{n}_{\mathrm{min}}T_{\mathrm{b}}+T_{\mathrm{l}})$ then
32             $[\xi_{\mathrm{opt}},n_{\mathrm{opt}}]\leftarrow[0,0]$
33             $\mathcal{D}.\mathrm{append}\left([t_{s},i_{s}]\rightarrow[\xi_{\mathrm{opt}},n_{\mathrm{opt}}]\right)$
34       else if $i_{s}==1$ then
35             $\left[\xi_{\mathrm{opt}},n_{\mathrm{opt}}\right]\leftarrow\left[\varepsilon_{s}(n_{\mathrm{opt}}),\left\lfloor\frac{t_{s}-T_{\mathrm{l}}}{T_{\mathrm{b}}}\right\rfloor\right]$
36             $\mathcal{D}.\mathrm{append}\left([t_{s},i_{s}]\rightarrow[\xi_{\mathrm{opt}},n_{\mathrm{opt}}]\right)$
37       else
38             $n_{\mathrm{opt}}\leftarrow\arg\max\limits_{n}\{\overline{\varepsilon}(n)\textnormal{{BestSucc!}}(t_{s}-nT_{\mathrm{b}}-T_{\mathrm{l}},i_{s},f_{G},\varepsilon_{\mathrm{max}},\mathcal{D})+$
39             ${}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}[1-\overline{\varepsilon}(n)]\textnormal{{BestSucc!}}(t_{s}-nT_{\mathrm{b}}-T_{\mathrm{l}},i_{s}-1,f_{G},\varepsilon_{\mathrm{max}},\mathcal{D})\}$
40             $\xi_{\mathrm{opt}}\leftarrow\overline{\varepsilon}(n)\textnormal{{BestSucc!}}(t_{s}-nT_{\mathrm{b}}-T_{\mathrm{l}},i_{s},f_{G},\varepsilon_{\mathrm{max}},\mathcal{D})+$
41             ${}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}[1-\overline{\varepsilon}(n)]\textnormal{{BestSucc!}}(t_{s}-nT_{\mathrm{b}}-T_{\mathrm{l}},i_{s}-1,f_{G},\varepsilon_{\mathrm{max}},\mathcal{D})$
42             $\mathcal{D}.\mathrm{append}\left([t_{s},i_{s}]\rightarrow[\xi_{\mathrm{opt}},n_{\mathrm{opt}}]\right)$
43       end if
44
45
46return $\xi_{\mathrm{opt}}$
47
48
49function ScheduleTreeGen($t_{s},i_{s},k_{s},\tau_{s},\mathcal{D}_{\mathrm{local}},T_{\mathrm{b}},T_{\mathrm{l}}$):
50       if $\mathcal{D}_{\mathrm{local}}.\mathrm{find}([t_{s},i_{s}])==\emptyset$ then
51             $\mathfrak{T}_{s}\leftarrow\emptyset$
52       else
53             $(\xi,n)\leftarrow\mathcal{D}_{\mathrm{local}}.\mathrm{find}([t_{s},i_{s}])$
54             $\mathfrak{T}_{s}\leftarrow\left(\textnormal{{ScheduleTreeGen}}(t_{s}-nT_{\mathrm{b}}-T_{\mathrm{l}},i_{s},k_{s}+1,\tau_{s}+nT_{\mathrm{b}},\mathcal{D}_{\mathrm{local}},T_{\mathrm{b}},T_{\mathrm{l}}),\{(t_{s},i_{s},k_{s},\tau_{s})\},\right.$
55             ${}~{}~{}~{}~{}~{}~{}~{}\left.\textnormal{{ScheduleTreeGen}}(t_{s}-nT_{\mathrm{b}}-T_{\mathrm{l}},i_{s}-1,0,0,\mathcal{D}_{\mathrm{local}},T_{\mathrm{b}},T_{\mathrm{l}})\right)$
56       end if
57
58
59return $\mathfrak{T}_{s}$
Algorithm 1 Integer DP for DMH-HARQ schedule optimization
Nevertheless, it shall be remarked that the real-value relaxation $n_{s}\in\mathbb{R}^{+}$ was taken, like in classical FBL works, only for the convenience of analysis. Indeed, representing the blocklength per transmission in the unit of channel use, the final solution $n_{\mathrm{opt}}(s)$ can only take integer values from $\mathbb{N}^{+}$. Meanwhile, to apply FBL approaches while guaranteeing to meet the stringent latency requirement, as referred earlier, $n_{s}$ is strictly constrained by (III-C). Thus, both the action set and the feasible region of status are usually of reasonable sizes, enabling us to apply classical integer DP techniques to directly solve the global optimum (the feasibility and optimality of this approach are discussed in [33]). A typical implementation is to recursively solve $n_{\mathrm{opt}}(s)$ from the last hop where $i_{s}=1$ on, hop-by-hop onto the first one where $i_{s}=I$. The recursive algorithm is essentially accompanied with LUT to store the optimal achievable utility over the solution space, as described by Algorithm 1:
  • Two global dictionaries, $\mathcal{D}_{\mathrm{local}}$ and $\mathcal{D}_{\mathrm{succ}}$, are defined in the main optimizing function OptSchedule as LUT s for the solutions of local hop (based on perfect CSI) and successor hops (based on statistical CSI), respectively. Each LUT maps the state space $\mathcal{S}$ onto $\mathbb{R}^{+}$ to record the achievable utility $\xi_{\mathrm{max}}(s)$.
  • The functions Best! and BestSucc! are implemented to recursively solve the optimal blocklength allocation policy $n_{s}$ and the associated best utility $\xi_{\mathrm{max}}(s)$ regarding perfect and imperfect CSI, respectively. Especially, BestSucc! is a deviated version of Best! that considers with the only statistical CSI for each hop, while Best! exploits the perfect CSI of the local hop. BestSucc! is called by Best! to estimate the achievable utility over successor hops. The global dictionaries dictionary $\mathcal{D}_{\mathrm{local}}$ and $\mathcal{D}_{\mathrm{succ}}$ are therewith updated, respectively.
  • The function ScheduleTreeGen is implemented to recursively generate the binary schedule tree $\mathfrak{T}$ from the dictionary $\mathcal{D}_{\mathrm{local}}$ that stores the blocklength allocation policy.
V-B Hop-by-Hop CSI Update and Reschedule
Note that by running Algorithm 1 at node $i\in\mathcal{I}$, the optimal schedule $\mathfrak{T}_{s}$ is solved based on the perfect CSI of the $i^{\mathrm{th}}$ hop and statistical CSI of all hops thereafter, which shall not be adopted by any other node. Indeed, upon a successful decoding of the message received from its predecessor node, each node $i\in\mathcal{I}$ shall individually solve its own optimal schedule $\mathfrak{T}_{s}$ regarding the instantaneous system state $s$ and its real-time channel measurement $g_{i}$ of the $i^{\mathrm{th}}$ hop. The complete procedure can be briefly summarized as in Algorithm 2.
1
2Initialization: $[g_{1},g_{2},\dots g_{I}],f_{G},\varepsilon_{\mathrm{max}},T_{\mathrm{max}},T_{\mathrm{b}},T_{\mathrm{l}}$
3 $s\leftarrow(T_{\mathrm{max}},I,0,0)$
4 for $i=I,I-1,\dots,1$ do
5       $\mathfrak{T}_{s}\leftarrow$OptSchedule($s,g_{i},f_{G},\varepsilon_{\mathrm{max}},T_{\mathrm{b}},T_{\mathrm{l}}$)
6       $s^{\prime}\leftarrow s$
7       while true do
8             (Re)-transmit regarding $\mathfrak{T}_{s}$
9             Update $s$ regarding the transmission result
10             if $n_{s}=0$ then
                   return
                    // Termination state
11
12             else if $i_{s}<i_{s^{\prime}}$ then
                   break
                    // Current hop succeeded
13
14             end if
15
16       end while
17
18 end for
return
Algorithm 2 $I$-hop DMH-HARQ with CSI update
V-C Computational Complexity Analysis
As described above, Algorithm 1 implements a recursive DP approach to compute the integer dynamic program. The recursive algorithm considers the DMH-HARQ as an $I$-stage DP problem, each stage representing a hop. Meanwhile, each hop-level sub-problem that starting with state $s$ makes also an $\mathcal{O}(n_{\mathrm{max},s})$-stage DP problem itself, each stage representing an available channel use in the schedule.
According to the computational complexity analysis of the CLARQ algorithm [33], which has the same structure as each hop-level sub-problem, we know that the time complexity of each hop-level sub-problem is $\mathcal{O}(n_{\mathrm{max},s})$. Since for each $s\in\mathcal{S}$ hop, $n_{\text{max},s}$ is upper-bounded by $n_{\mathrm{MAX}}\triangleq\frac{t_{s}-T_{\mathrm{min}}^{(I-1)}}{Mf_{\mathrm{s}}}$, the time complexity of solving the optimal DMH-HARQ at the $i^{\mathrm{th}}$ last node is ${\mathcal{O}\left(n_{\mathrm{MAX}}^{i}\right)}$. Taking the hop-by-hop execution of Algorithm 2 into account, the overall time complexity of the DMH-HARQ approach over the $I$-hop relay chain is $\mathcal{O}\left[\sum\limits_{i=1}^{I}n_{\mathrm{MAX}}^{i}\right]=\mathcal{O}\left(n_{\mathrm{MAX}}^{I}\right)$.
In realistic scenarios of deployment, such a time complexity can critically challenge the online computation of optimal DMH-HARQ schedule, especially due to the time-varying channel gains $g_{i}$, which leads to high delay and significant power consumption. To address this issue, it is a practical solution to implement a LUT in each node, which that contains a set of offline solved dictionaries regarding various channel conditions. Since the operation of searching for a specific entry in a dictionary-based LUT has only a time complexity of $\mathcal{O}(1)$, the devices can be rapidly adapted to the appropriate specification w.r.t. the real-time channel measurement, leading to an overall time complexity as low as $\mathcal{O}(I)$ for the entire $I$-hop DMH-HARQ process.
VI Numerical Evaluation
To evaluate our proposed approaches, we conducted numerical simulations to benchmark DMH-HARQ against baseline solutions with respect to different specifications of the system and the environment.
TABLE I: Setup of the simulation campaign
Parameter Value Remark
[…]
$M$ $2$ Modulation order, 2 for BPSK
[…]
$\varepsilon_{\mathrm{max}}$ $0.5$ Maximal allowed PER per transmission
[…]
$\mathbf{T}_{\mathrm{max}}$$\{0.4,0.5,\dots,1\}$ $\mathrm{ms}$ Test range of time frame length
$I$ $4$ Default number of hops
$\mathbf{I}$$\{2,3,\dots,10\}$ Test range of hops number
$L_{\text{f},i}$ $\sim\mathrm{Rice}(0.5,1)$ Small-scale fading with a consistent LoS path
[…]
$\overline{\mathbf{g}}$$\{-10,-9,\dots,0\}$ $\mathrm{dB}$ Test range of mean channel gain
$n_{\mathrm{MC}}$ $1~{}000~{}000$ Runs per Monte-Carlo test
VI-A Baseline Methods
Three baseline methods are implemented and evaluated in our simulations:
  1. 1.
    Naive static HARQ -IR: where the (re)transmission times on each hop and coding rate of every individual (re)transmission are pre-scheduled. For each hop, a full IR (Type II HARQ) is applied to combine the codes sent in different transmissions. Since the performance is highly dependent on the (re)transmission times and the coding rate, here we consider the approximate bound provided by Makki et al. in [5], which is only achievable with ideal Gaussian codebooks allowing non-integer blocklength. Under a naive schedule, the same length of time is assigned to every hop. Recall that the performance of Type I HARQ is guaranteed outperformed by that of Type II, and that our DMH-HARQ approach is evaluated with Type I HARQ. Therefore, the benchmark is unfairly favoring the baseline over our approach.
  2. 2.
    Optimal static HARQ-IR: which follows the same principle of pre-scheduled retransmission as the naive static HARQ-IR. However, the sub-frame lengths allocated to different hops are jointly optimized so as to minimize the E2E packet loss rate. Again, Type II HARQ-IR is applied and the bound in [5] is considered, granting this baseline an unfair advantage in competition against our approach.
  3. 3.
    Listening-based cooperative ARQ: which was proposed by Goel et al. in [36]. Similar to the naive static HARQ approach, every hop is assigned with the same length of sub-frame and pre-scheduled with the same number of (re)transmissions and coding rate. However, if the DF on some hop $i$ succeeds before using up all the pre-scheduled (re)transmission slots assigned to it, the DF on hop $i+1$ will immediately be triggered, inheriting all the remaining time resource from its predecessor hop. In our simulation, each hop is pre-scheduled with $2$ (re)transmission slots, applying Type-I HARQ without combining.
The principles of both static HARQ-IR and listening-based cooperative ARQ are illustrated in Fig. 2 for comparison with Fig. 1. Note that the naive and optimal schemes of static HARQ-IR only distinguish from each other by weather the sub-frames are of of same length over all hops.
Figure 2: Illustration of baseline solutions for 4-hop DF chain.
VI-B Simulation Setup
The setup of our simulation campaign is detailed in Tab. I. Among all the parameters, there are four that are of particular interest to us to test the sensitivity of our approach, namely the total time frame length $T_{\mathrm{max}}$, the number of hops $I$, the decoding/feedback delay $T_{\mathrm{l}}$, and the mean channel gain $\overline{g}$. Each of these key parameters is assigned with a default value and a test range. When benchmarking the approaches regarding one key parameter over its test range, all the rest key parameters are specified to their default values. For every individual specification, $n_{\mathrm{MC}}=1~{}000~{}000$ runs of Monte-Carlo test are conducted, each run with an independently generated set of channel conditions for the $I$ hops, with which all approaches are tested.
VI-C Results and Analyses
The results of sensitivity tests regarding $T_{\mathrm{max}}$, $I$, $T_{\mathrm{l}}$ and $\overline{g}$ are illustrated in Fig. 3. Generally, the E2E packet loss rate monotone increase as i) the average blocklength that can be allocated to each hop (exclusive the decoding/feedback delay) decreases; and ii) the SNR decreases.
We observe from the results that the DMH-HARQ approach consistently outperforms all baselines, even when under an unfair disadvantage in soft combining. It is promised to hold even more performance gain over the static HARQ solutions when applied with a more advanced combining technology. This excellent performance is credited to the DP algorithm that guarantees to achieve the theoretical maximum of the Bellman sum (12). Especially, with more available blocklength per hop, since the overall possible numbers of HARQ attempts over the $I$-hop chain increases, this performance gain brought by DP is also increasing. To the contrary, when the channel is harsh or the radio resource is extremely limited, the constraint (III-C) is likely forcing to allow only one transmission slot per hop, so that the optimal DMH-HARQ schedule converges to the optimal one-shot-per-hop schedule, with just gives the performance upper bound of optimal static HARQ-IR in FBL [5].
Figure 3: Sensitivity tests regarding (3(a)) the time frame length, (3(b)) the number of hops, (3(c)) the decoding/feedback delay, and (3(d)) the pathloss, respectively.
Remark that the listening-based cooperative ARQ approach also allocates the radio resource to different hops in a technically dynamic manner. However, its mechanism design only allows the successor hops to inherit unused resource from predecessors, not the other way around. With a same fixed blocklength for every (re)transmission, it results in a longer effective sub-frame for the late hops that are closer to the final information sink than the for early hops that are closer to the source. However, as it can be derived from [37], as a stochastic DP, the optimal solution shall be just the opposite: to allocate more resource to the early hops than the late ones. In the end, the listening-based cooperative ARQ approach cannot hit the optimum of the DP problem (III-C), and has only a limited gain from the dynamic scheduling. This impact is even more critical in the FBL regime where ARQ/HARQ generally work poor, leading to a significant gap to the performance in the infinite blocklength (IBL) regime. As we see from the results, the listenting-based cooperative ARQ approach performs the worst among all baselines.
VII Further Discussions
VII-A Computational Complexity
Concerning the recursive nature of Algorithm 1, its computational load can be a crucial factor in assessing its practical applicability. Since DMH-HARQ can be considered as an $I$-hop generalization of the $2$-hop CLARQ protocol, we can follow the approach we used in [33] to analyze its time complexity, which returns us a result of $\mathcal{O}\left(n_{\mathrm{MAX}}^{I}\right)$.
It must be noted that, though this time complexity is significantly higher than to optimize the schedule of conventional static HARQ approaches, i.e. $\mathcal{O}\left[\log\left(n_{\mathrm{MAX}}\right)\right]$, the optimizer is not necessarily executed in real time for every communication session. Indeed, the optimal schedules regarding different CSI can be obtained offline for every node with Algorithm 1, and cumulatively stored in a LUT. In the runtime, the node only needs to look up the table for its optimal decision, which has a minimal time complexity of $\mathcal{O}\left(1\right)$.
VII-B IR Gain
We have proven that even when applied without any combining technique, DMH-HARQ is able to outperform the optimal static HARQ-IR. A natural idea, of course, is to further enhance the performance of DMH-HARQ with soft combining, especially with an ideal IR, where the PER formula (8) must be replaced by
$\varepsilon_{s}=Q\left(\sqrt{\frac{n_{s}+\frac{\tau_{s}}{T_{\mathrm{b}}}}{V_{I+1-i_{s}}}}\left(\mathcal{C}_{I+1-i_{s}}-\frac{d}{n_{s}+\frac{\tau_{s}}{T_{\mathrm{b}}}}\right)\ln 2\right).$ (14)
Though IR is guaranteed to reduce the achievable E2E packet loss rate of DMH-HARQ, it costs more computational effort. The inclusion of $\tau_{s}$ into $\varepsilon_{s}$ rejects Lemma 1. To keep it an MDP so that the integer DP based framework of Algorithm 1 still applies, $\tau_{s}$ must play its role as part of the system state when solving the decision policy, i.e. $n_{s}$. The dictionaries $\mathcal{D}_{\mathrm{local}}$ and $\mathcal{D}_{\mathrm{succ}}$ in Algorithm 1 shall map $[t_{s},i_{s},\tau_{s}]$ instead of $[t_{s},i_{s}]$ to $[\xi_{\mathrm{opt}},n_{\mathrm{opt}}]$. This will significantly increase the effective size of the state space, raising the time complexity of solving $n_{s}$ at each node from $\mathcal{O}(n_{\mathrm{MAX}}^{I})$ to $\mathcal{O}(n_{\mathrm{MAX}}^{2I})$.
VII-C Multi-Access Design
In this paper we have been focusing on a single multi-hop data link. In practical deployment, a multi-access solution will be mandatory to allow efficient networking, especially in the fronthaul and midhaul domains, where the network exhibits a star topology that each DU is associated with multiple RUs and each CU with multiple DUs. While OFDMA is widely used in modern radio systems and has been dominating most wireless standards, it may not be the optimal match for our DMH-HARQ protocol, due to the challenge in interference management caused by the dynamic traffic pattern that are hard to predict. Similar issue has already been discussed for the original CLARQ problem [33], revealing that time-division multiple access (TDMA) will be a better solution than OFDMA in this context. This will limit the application of the dynamic HARQ approaches such like CLARQ or DMH-HARQ for the air interface between users and base stations, which are generally standardized in modern cellular systems to use OFDMA for better scalability and flexibility. It is also for this reason that we exclude the air interface from the scope of this study.
However, in wireless transport networks which is under our focus in this study, TDMA can still be a good and feasible option, because: i) unlike the air interface of Radio Access Network (RAN), the physical layer design of transport network is not universally standardized but upon the design of vendor, allowing to use TDMA; ii) compared to the number of UEs in RAN, the number of transport network nodes are limited, leading to low demand for scalability; iii) the transport network nodes are generally installed at fixed position, leading to easy synchronization and low demand for flexibility.
Moreover, the TDMA design in DMH-HARQ can be potentially combined with the target wake time (TWT) mechanism, which is utilized by the IEEE 802.11ax standard, to further raise the time efficiency of DMH-HARQ: after successfully forwarding a message, the relaying node can send a triggering token to the next transmitting node, so that it is immediately waken from sleep and start its the DMH-HARQ process, instead of waiting till the last transmitting node completing its multi-hop relay. In such way, multiple data links can be parallelized in a pipeline fashion over the multiple hops, so that the E2E delay is reduced.
VIII Conclusion and Outlooks
In this paper, we have looked into the problem of dynamic HARQ in multi-hop wireless transport network with a special focus on reliability and openness. Considering FBL regime, we have proposed a DMH-HARQ scheme with an associated DP algorithm that optimizes it. The proposed methods are proven via numerical simulations as effective and outperforming conventional baselines, even without any code combining. Its performance superiority remains consistent regarding the link delay budget and decoding delay, and increases along with the hop number and the channel quality.
As for the next step, we plan to apply IR to DMH-HARQ, and extend our analyses and optimization methods in this paper thereto. Deeper study to the multiple access design for DMH-HARQ and its integration with TWT mechanism is also an interesting topic for future study.
Acknowledgment
This work is partially supported by the German Federal Ministry of Education and Research in the programme of “Souverän. Digital. Vernetzt.” joint projects 6G-RIC (16KISK028) and Open6GHub (16KISK003K/16KISK004/16KISK012). This work is partially supported by SNS JU Project 6G-GOALS (GA no. 101139232).
Appendix A Proof of Lemma 2
Proof.
For the last hop ($i_{s}=1$), consider an arbitrary DMH-HARQ scheme with maximal $K$ transmission attempts, where the blocklength of $k^{\mathrm{th}}$ (re-)transmission is $m_{k}$ for all $k\in\{1,2,\dots K\}$. Obviously, since $i_{s}=1$, the utility equals the overall chance of successful transmission. Using Type I HARQ without combining, that is
$\begin{split}\xi=&1-\epsilon_{I}(m_{1})+\epsilon_{I}(m_{1})\left[1-\epsilon_{I}(m_{2})\right]\\ &+\epsilon_{I}(m_{1})\epsilon_{I}(m_{2})\left[1-\epsilon_{I}(m_{3})\right]+\dots\\ &+\epsilon_{I}(m_{1})\dots\epsilon_{I}(m_{K-1})\left[1-\epsilon_{I}(m_{K})\right]\\ =&1-\prod\limits_{k=1}^{K}\epsilon_{I}(m_{k})\end{split}$ (15)
Meanwhile, according to [5], if it is Type-II HARQ with IR applied, the overall chance of successful transmission is equivalent to that of an one shot transmission:
$\xi_{\mathrm{os}}=1-\epsilon_{I}\left(\sum\limits_{k=1}^{K}m_{k}\right).$ (16)
Obviously, under the same blocklength allocation and channel conditions, Type II HARQ-IR always outperforms Type I HARQ without combining in error rate, i.e., $\prod\limits_{k=1}^{K}\epsilon_{I}(m_{k})\geqslant\epsilon_{I}\left(\sum\limits_{k=1}^{K}m_{k}\right)$, which takes the equality only when $K=1$. Furthermore, given the fixed remaining time $t_{s}$, we always have the constraint $\sum\limits_{k=1}^{K}\left(m_{k}T_{\mathrm{b}}+T_{\mathrm{l}}\right)\leqslant t_{s}$, which means
$\sum\limits_{k=1}^{K}m_{k}\leqslant\frac{t_{s}-KT_{\mathrm{l}}}{T_{\mathrm{b}}}\leqslant\frac{t_{s}-T_{\mathrm{l}}}{T_{\mathrm{b}}},$ (17)
where the first equality is taken only when $t_{s}$ is sufficiently exploited, and the second taken only when $K=1$. Additionally, i always holds that $\partial\epsilon_{I}(m)/\partial m<0$ for all $m\in\mathbb{R}^{+}$. Thus, we have
$\prod\limits_{k=1}^{K}\epsilon_{I}(m)\geqslant\epsilon_{I}\left(\frac{t_{s}-T_{\mathrm{l}}}{T_{\mathrm{b}}}\right),$ (18)
where the equality is only taken when $K=1$ and $m_{1}=\frac{t_{s}-T_{\mathrm{l}}}{T_{\mathrm{b}}}$, and the Lemma is therewith proven.
Appendix B Proof of Theorem 1
Proof.
For the first step, we prove the monotone of $\xi_{\mathrm{max}}(s)$ regarding $t_{s}$ as follows. Given any feasible system status $s_{0}=(t_{s_{0}},i_{s_{0}},k_{s_{0}},\tau_{s_{0}})$, let $n_{\mathrm{opt}}$ be an optimal blocklength allocation policy [^1] The existence of such optima is guaranteed by the bounded range of $\xi(s_{0})$. However, the uniqueness of $n_{\mathrm{opt}}$ is not guaranteed. Hence we use the term “an optimal policy” instead of “the optimal policy” here. that generates the DMH-HARQ schedule $\mathfrak{T}_{s_{0}}^{\mathrm{opt}}$ with maximized utility $\xi_{\mathrm{max}}(s_{0})$. According to Lemma 2, there must be $n_{\mathrm{opt}}(s)=\frac{t_{s}}{T_{\mathrm{b}}}-T_{\mathrm{l}}$ for all $s$ that $i_{s}=1$.
Given an arbitrary $\Delta t>0$, let $s^{<\Delta t>}$ denote $(t_{s}+\Delta t,i_{s},k_{s},\tau_{s})$, we can always construct a sub-optimal $n_{\mathrm{sub}}$ that
$n_{\mathrm{sub}}(s)=\begin{cases}n_{\mathrm{opt}}(s^{<-\Delta t>})+\frac{\Delta t}{T_{\mathrm{b}}}&i_{s}=1,\\ n_{\mathrm{opt}}(s^{<-\Delta t>})&\mathrm{otherwise}.\end{cases}$ (19)
Now apply $n_{\mathrm{sub}}$ on the initial state $s_{1}=s_{0}^{\Delta t}$ to generate another DMH-HARQ schedule $\mathfrak{T}^{\mathrm{sub}}_{s_{1}}$, it is straightforward to derive that $\mathfrak{T}^{\mathrm{opt}}_{s_{0}}$ and $\mathfrak{T}^{\mathrm{opt}}_{s_{1}}$ have the identical structure, except that for each non-empty node $s$ in $\mathfrak{T}^{\mathrm{opt}}_{s_{0}}$, the corresponding node of $\mathfrak{T}^{\mathrm{opt}}_{s_{1}}$ is
$s^{\prime}=\begin{cases}s&i_{s^{\prime}}=0,\\ s^{<\Delta t>}&\text{otherwise}.\end{cases}$ (20)
And error rates of each transmission attempt regarding the two schedules, denoted as $\varepsilon_{s}$ and $\varepsilon^{\prime}_{s^{\prime}}$, respectively, must fulfill
$\varepsilon^{\prime}_{s^{\prime}}\begin{cases}<\varepsilon_{s}&i_{s^{\prime}}=1,\\ =\varepsilon_{s}&\text{otherwise}.\end{cases}$ (21)
Recalling Eq. (12), it always hold $\xi(\mathfrak{T}^{\mathrm{sub}}_{s_{1}})>\xi_{\mathrm{max}}(s_{0})$. Meanwhile, there is always some $n^{\prime}_{\mathrm{opt}}$ that generates an optimal schedule $\mathfrak{T}^{\mathrm{opt}}_{s_{1}}$ with maximal $\xi_{\mathrm{max}}(s_{1})\geqslant\xi(\mathfrak{T}^{\mathrm{sub}}_{s_{1}})$. Thus:
$\xi_{\mathrm{max}}(s_{1})=\xi_{\mathrm{max}}(s_{0}^{<\Delta t>})>\xi_{\mathrm{max}}(s_{0})\quad\forall s_{0},\forall\Delta t>0,$ (22)
i.e. $\xi_{\mathrm{max}}(s)$ is strictly monotone increasing regarding $t_{s}$.
Next, we come to prove the continuity of this function. According to Eq. (8), $\varepsilon_{s}$ is continuous function of $t_{s}$, so $\xi_{\mathrm{max}}(s)$ can be guaranteed $t_{s}$-continuous, if both $\xi_{\mathrm{max}}(s_{\mathrm{l}}(\mathfrak{T}_{s}))$ and $\xi_{\mathrm{max}}(s_{\mathrm{r}}(\mathfrak{T}_{s}))$ are continuous functions of $t_{s}$. This sufficient condition can be derived via mathematical induction as follows.
Starting with the two termination cases of $t_{s}<T_{\mathrm{min}}^{(i_{s})}$ and $i_{s}=0$, respectively, according to Eqs. (1) and (2) we have $S(L(\mathfrak{T}_{s}))=S(R(\mathfrak{T}_{s}))=\emptyset$. Recalling (5), for the termination cases we have $\xi_{\mathrm{max}}(s)=\xi(\mathfrak{T}_{s})\equiv 0$, which are both $t_{s}$-continuous. Therefore, the utility $\xi_{\mathrm{max}}(s)$ itself is also $t_{s}$-continuous.
Then consider the last hop where $i_{s}=1$, according to Lemma 2 we know that $\xi_{\mathrm{max}}(s)$ is achieved when $n_{s}=\frac{t_{s}-T_{\mathrm{l}}}{T_{\mathrm{b}}}$, leading to a left branch where $t_{s_{\mathrm{l}}(\mathfrak{T}_{s})}=0<T_{\mathrm{min}}^{(1)}$ and a right branch where $i_{s_{\mathrm{r}}(\mathfrak{T}_{s})}=0$, both are termination cases. Thus, as proven above we have both $\xi_{\mathrm{max}}(s_{\mathrm{l}}(\mathfrak{T}_{s}))$ and $\xi_{\mathrm{max}}(s_{\mathrm{r}}(\mathfrak{T}_{s}))$ are $t_{s}$-continuous, and therefore $\xi_{\mathrm{max}}(s)$ the same as well.
For other hops other than the last one, i.e. $i_{s}\in\{2,3,\dots I\}$, according to the DMH-HARQ principle (1) and (2), it is obvious that there are only three possible combinations of its left and right branches:
  1. 1.
    $s_{\mathrm{l}}(\mathfrak{T}(s))=\emptyset,i_{s_{\mathrm{r}}(\mathfrak{T}(s))}=\emptyset$;
  2. 2.
    $s_{\mathrm{l}}(\mathfrak{T}(s))=\emptyset,i_{s_{\mathrm{r}}(\mathfrak{T}(s))}=i_{s}-1$;
  3. 3.
    $i_{s_{\mathrm{l}}(\mathfrak{T}(s))}=i_{s},i_{s_{\mathrm{r}}(\mathfrak{T}(s))}=i_{s}-1$.
For case 1), $\xi_{\mathrm{max}}(s)\equiv 0$, which is $t_{s}$-continuous. For case 2), $\xi_{\mathrm{max}}(s_{\mathrm{l}}(\mathfrak{T}_{s}))$ is $t_{s}$-continuous as proven above, so that $\xi_{\mathrm{max}}(s)$ can be guaranteed $t_{s}$-continuous if only $\xi_{\mathrm{max}}(s_{\mathrm{r}}(\mathfrak{T}_{s}))$ is $t_{s}$-continuous. For case 3), we recall Lemma 1 that $n_{\mathrm{opt}}(s)$ depends only on $t_{s}$ and $i_{s}$, so that $\xi_{\mathrm{max}}(\mathfrak{T}(s))$ and $\xi_{\mathrm{max}}(L(\mathfrak{T}(s)))$ are different samples for the same function of $t_{s}$. Since the iterative generation of the left branch in decision tree is determined to terminate with an empty set, i.e. case 1) or 2), we can conclude that for a $\mathfrak{T}s$ in case 3), both $\xi_{\mathrm{max}}(s)$ and $\xi_{\mathrm{max}}(s_{\mathrm{l}}(\mathfrak{T}(s)))$ are guaranteed $t_{s}$-continuous if $\xi_{\mathrm{max}}(s_{\mathrm{r}}(\mathfrak{T}_{s}))$ is $t_{s}$-continuous. Since the iterative generation of the right branch in decision tree always ends up with one of the two termination cases, which are both $t_{s}$-continuous, we can conclude from the analyses above that $\xi_{\mathrm{max}}(s)$ is also $t_{s}$-continuous for all $s$ where $i_{s}>1$.
Thus, the continuity is proven for all $i\in\mathcal{S}$. With both the monotone and the continuity derived, the theorem is proven. ∎